Thomas’ Calculus Early Transcendentals 13ed

Homework Assignment for Chapter 06

1. Section 6.1: problems 5, 23, 55, 63.

2. Section 6.2: problem 1, 27, 39.

3. Section 6.3: problems 2, 9, 25, 29.

4. Section 6.4: problem 13, 19, 23.